OpenStudy (anonymous):

Caleb is studying for his end-of-year exams. The graph below shows the number of hours he studied for the first six days. A line of best fit has been drawn on the graph. Scatter plot with line of best fit in the first quadrant titled Calebs Studying Time, shows Day on the x axis and Number of Hours Studied on the y axis. The x axis scale is from 0 to 8 in increments of 1. The y axis scale is from 0 to 7 in increments of 1. The line of best fit passes through the points one comma one and four comma two. Based on the line of best fit, for how many hours will Caleb study on day 10? 2 3 4 5

OpenStudy (whpalmer4):

I can't see the graph :-)

OpenStudy (anonymous):

Is it four?

OpenStudy (whpalmer4):

well, let's work out the equation: slope = \(m = \frac{y_2-y_1}{x_2-x_1} = \frac{2-1}{4-1} = 1/3\) We know the line passes through the point (1,1) so we can write with the point-slope formula \[(y-y_0 = m(x-x_0) \]\[y-2= \frac{1}{3}(x-4)\]Plug in x = 10 \[y - 2 = \frac{1}{3}(10-4)\]\[y = 2 + 2 = 4\]

OpenStudy (anonymous):

You re the bestest. Thanks

OpenStudy (whpalmer4):

of course, any decent piece of software that had done the fit of the line for you would have also given you flippin' equation, but no, not here :-)